OpenStudy (anonymous):

HELP!!!! -12 divided by 20/21

OpenStudy (emmynimmy):

-0.0285

OpenStudy (anonymous):

thx :)

OpenStudy (anonymous):

wait just to let u know 20/21 is a fraction

OpenStudy (anonymous):

@emmynimmy

OpenStudy (emmynimmy):

ohhhh

OpenStudy (emmynimmy):

i thought it was divided.

OpenStudy (emmynimmy):

im not sure I'm so sorry!!!!

OpenStudy (anonymous):

ok but u can keep the medal

OpenStudy (emmynimmy):

Hahah, thank you.

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@slipknot306

OpenStudy (slipknot306):

what grade u in

OpenStudy (anonymous):

7th

OpenStudy (capnkookcx):

= -12/1 ÷ 20/21 = (-12 × 21) / (20 × 1) = -252/20 = -63/5 = -12 3/5

OpenStudy (anonymous):

:/

OpenStudy (slipknot306):

k so -12u3/5

OpenStudy (anonymous):

Consider the following fractions.

OpenStudy (anonymous):

7/6, 8/3, 9/4, 11/9

OpenStudy (anonymous):

Determine which of the fractions is a terminating decimal number. Plot the point representing that fraction on a number line.

OpenStudy (slipknot306):

o um look im not really good with fractions i barely made it to 8th grade

OpenStudy (anonymous):

ok :(

OpenStudy (slipknot306):

but @Capnkookcx

OpenStudy (anonymous):

can u plz help me

OpenStudy (capnkookcx):

Give me a second

OpenStudy (slipknot306):

@Miracrown

OpenStudy (anonymous):

@OregonDuck

OpenStudy (anonymous):

@YoungStudier

OpenStudy (capnkookcx):

7/6 = 1.166666666666667 8/3 = 2.666666666666667 9/4 = 2.25 11/9 = 1.222222222222222 So I would say 2.25

OpenStudy (youngstudier):

@princeevee

OpenStudy (anonymous):

@Capnkookcx

OpenStudy (anonymous):

@Capnkookcx

OpenStudy (anonymous):

never mind :(

OpenStudy (slipknot306):

yo weggy will help you

OpenStudy (slipknot306):

weggy.com